OpenStudy (anonymous):

I don't get it!!! Review the steps used to solve the problem below. Explain whether the steps are correct or incorrect. If you find an incorrect step, explain what you would do to fix the mistake. 4 • (x + 2) = 14 Step 1: 4x + 2 = 14 Step 2: 4x = 12 Step 3: x = 3

OpenStudy (aripotta):

step 1 is incorrect. the 4 was distributed, but only the x. it should have also been distributed to the 2. it should have looked like: 4x + 8 = 14. then subtract 8 from both sides to get: 4x = 6. and then divide 4 from both sides to get x = 1.5

OpenStudy (anonymous):

I still don't understand.

OpenStudy (anonymous):

(x+2)=14 Remove the parentheses around the expression x+2. x+2=14 Since 2 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 2 from both sides. x=-2+14 Add 14 to -2 to get 12. x=12

OpenStudy (aripotta):

it was 4 x (x + 2) = 14